A solution has a pH of 4.57. Calculate the pOH, calculate [H₃O⁺], and calculate [OH⁻].
1
Expert's answer
2020-05-13T14:20:49-0400

pOH = 14 - pH = 14 - 4.57 = 9.43

[H3O+] = 10^-pH = 10^(-4.57) = 2.69*10^-5

[OH-] = 10^-pOH = 10^(-9.43) = 3.72*10^-10